A satellite is in a circular orbit around the Earth. It has a total energy (K + U) of -8.9x1010 J. Determine the Kinetic energy for this satellite in this orbit.

  1. 0.8x10^10 J
  2. 1.3x10^10 J
  3. 4.5x10^10 J
  4. 8.9x10^10 J

10.

How much energy would be needed to boost the satellite in question #9 to an infinite distance from the planet around which it is orbiting?

  1. 0.8x10^10 J
  2. 1.3x10^10 J
  3. 4.5x10^10 J
  4. 8.9x10^10 J
Answer #1

9.)

K.E = - total energy

Hence K.E = 8.9×10^10joule

10.)

Energy required to move to infinite from initial position = - Potential energy

=) -P.E = - total energy/2 = 4.5×10^10 Joule